### v1

# Tag&Tab: Pretraining Data Detection in Large Language Models Using Keyword-Based Membership Inference Attack

Abstract
Large language models (LLMs) have become essential digital task assistance tools. Their training relies heavily on the collection of vast amounts of data, which may include copyright-protected or sensitive information. Recent studies on the detection of pretraining data in LLMs have primarily focused on sentence-level or paragraph-level membership inference attacks (MIAs), usually involving probability analysis of the target model prediction tokens. However, the proposed methods often demonstrate poor performance, specifically in terms of accuracy, failing to account for the semantic importance of textual content and word significance. To address these shortcomings, we propose Tag&Tab, a novel approach for detecting data that has been used as part of the LLM pretraining. Our method leverages advanced natural language processing (NLP) techniques to tag keywords in the input text - a process we term Tagging. Then, the LLM is used to obtain the probabilities of these keywords and calculate their average log-likelihood to determine input text membership, a process we refer to as Tabbing. Our experiments on three benchmark datasets (BookMIA, MIMIR, and the Pile) and several open-source LLMs of varying sizes demonstrate an average increase in the AUC scores ranging from 4.1% to 12.1% over state-of-the-art methods. Tag&Tab not only sets a new standard for data leakage detection in LLMs, but its outstanding performance is a testament to the importance of words in MIAs on LLMs.
1 Introduction
The rapid advancement of generative artificial intelligence (GenAI) in recent years has significantly shifted the tech industry’s focus toward the development of powerful tools such as large language models (LLMs). Used for applications ranging from conversational chats to content generation [1, 2], LLMs have become an essential tool in the digital landscape. Notable examples of foundational LLMs and their applications include OpenAI’s GPT-4 [3], which is used in applications such as ChatGPT [4], Google’s Gemini [5], which is integrated in various Google services for enhanced user interactions and AI-driven features [6], and Meta’s LLaMA [7], which is used in both research and commercial applications for diverse language understanding and content generation tasks.
The rapid adoption and widespread use of these models in fields ranging from customer service to scientific research highlight the intense competition among tech companies and research institutions to develop the most advanced LLMs. While this competition drives performance improvements, it is heavily dependent on the collection of large amounts of data [8]. To achieve these improvements, LLMs are primarily trained on open-source datasets obtained from various sources using methods such as synthetic data generation and web scraping [9, 10]. The type of data collected may include books, code, academic papers, and medical records [11, 12, 13, 7]. The methods employed for data collection raise significant privacy and ethical concerns [14, 15], primarily regarding the inclusion of personally identifiable information (PII) [16] and copyright-protected content, whose inclusion is illegal, violating privacy and intellectual property rights [17, 18, 11].
Awareness regarding these privacy and ethical issues has increased as a result of legal conflicts and the lack of transparency regarding the data collection process [13, 19, 20]. The lawsuit between The New York Times and OpenAI [21], as well as other lawsuits [22, 23], highlight the critical need for mechanisms aimed at detecting such violations of privacy and intellectual property, and more specifically, identifying the data used to train LLMs [24]. Such detection can be challenging, and several techniques have been proposed for detecting LLMs’ pretraining data. Methods such as sentence-level membership inference attacks (MIAs) and paragraph-level MIAs are designed to determine whether a specific text was part of a model’s training dataset [25, 26].
To determine whether a specific piece of data was included in the training set, MIAs exploit the differences in a model’s behavior when it processes seen and unseen data. The underlying assumption is that different patterns in the LLM will be seen (e.g., higher prediction confidence or greater loss reduction) when a model encounters data it has seen during training compared to new data. Despite MIAs’ demonstrated effective performance, they have several flaws [27]. First, their performance in terms of common metrics, such as accuracy and the true positive rate (TPR) at a fixed low false positive rate (FPR), tends to worsen as the training set size increases, often rendering it close to random [28]. This is due to the tradeoff between generalization and memorization in LLMs: as the model is trained on more data, it will generalize better, while increasing the number of model parameters increases the model’s tendency to memorize the pretraining data [29, 30]. The second flaw of existing MIA methods is their reliance solely on calculating token probabilities for sections (or all) of a text. This approach fails to consider the contextual and semantic aspects of the text, which contains crucial information about text membership [31, 32]. Third, most MIAs exhibit inconsistent performance, excelling on certain models and datasets while failing to generalize across different models and data distributions [27, 24]. Lastly, MIAs are often evaluated on data that is not independently and identically distributed (IID), which can lead to the detection of distribution shifts rather than genuine membership inference, thereby undermining the attacks’ reliability [33].
To address the limitations of existing methods, we introduce Tag&Tab, a novel approach based on common natural language processing (NLP) methods which is designed to efficiently and effectively detect LLMs’ pretraining data. Specifically, our method aims to determine whether an LLM was trained on a given text sample, given black-box access to the target LLM (i.e., can only query the model). Similar to other previous studies, we also assume partial knowledge of the pretraining data, i.e., word distribution [25, 34].
Inspired by the research of Lukas et al. [16], our Tag&Tab method leverages common NLP techniques for keyword selection and their subsequent prediction by the inspected LLM. Their work demonstrates the importance of focusing on named entity recognition (NER) words in identifying PII leakage, an insight which influenced the design of our method and specifically the priority it places on selecting high-entropy words in a given text. Our method consists of the following steps:
  1. 1.
    Preprocessing - Constructing a word entropy map and filtering certain sentences to ensure optimal keyword selection.
  2. 2.
    Tagging - Identifying the high-entropy words in the text and selecting the $K$ words with the highest entropy value, referred to as keywords.
  3. 3.
    Tabbing - Passing the entire text to the target LLM and calculating the average log-likelihood of these $K$ keywords.
  4. 4.
    Inference - Comparing the average log-likelihood to a threshold to determine the text’s membership (i.e., whether it was in the inspected model training set).
The intuition behind our method is that a higher log-likelihood for these challenging-to-predict high-entropy keywords indicates that the model previously encountered the text during its training phase [29]. Based on this intuition, we hypothesize that these rare keywords are more likely to be memorized by the model and thus serve as effective indicators of the text’s membership in the pretraining dataset [35, 28]. Furthermore, by selecting just a small number of high-entropy keywords, our method captures the most informative parts of the text, minimizing noise from other words’ probabilities.
We evaluated our method on nine LLMs of varying sizes and across three datasets containing 8 types of textual data. Our results demonstrate Tag&Tab’s superior performance over the examined state-of-the-art (SOTA) MIAs, with our method achieving an AUC score improvement of over 10% compared to the best-performing SOTA attack across multiple types of textual data.
The contributions of our paper are as follows.
  • We propose Tag&Tab, a novel approach for the detection of LLMs’ pretraining data that focuses on the contextual and semantic relevance of the words in a text, and opens the door to additional research on MIAs against LLMs.
  • To the best of our knowledge, we are the first to create a robust reference-free MIA method that demonstrates high and consistent performance across multiple textual data types and LLMs, while outperforming SOTA methods.
  • Our proposed approach proves to be both resource and time efficient, unlike reference-based attacks, which necessitate training a reference model, and other reference-free attacks that require an additional model [34].
Figure 1: Illustration of the Tag&Tab method - The process starts by inputting a text (in this example, the input is the conclusion of the well-known poem “The Road Not Taken" [36]) in the target LLM to obtain its word probability distribution (word probability). In the tag step, the keywords are selected based on the words’ entropy value (created in the preprocessing phase). In the tab step, the log-likelihood of the selected keywords is calculated. Finally, in the infer step, the average log-likelihood of the chosen keywords is compared against a threshold $\gamma$ to determine if the text was part of the target LLM’s pretraining data.
2 Related Work
Membership Inference (MI) [37] is a classification task that aims to determine whether a data sample $x$ was part of the training data $D_{\text{train}}$ of a model $f$. The attacker receives a sample $x$ and a model $f$, and uses the attack model $A$ to classify $x$ as a member $x\in D_{\text{train}}$ if $A(f(x))=1$; otherwise, $x$ is classified as a non-member $x\notin D_{\text{train}}$.
Large Language Model Membership Inference is a subdomain of membership inference that has received increasing research attention. Within this subdomain, the topic of detecting pretraining data has been the focus of numerous studies exploring various methodologies for identifying whether specific texts were included in the training dataset of an LLM. Existing MIAs against LLMs can be categorized as reference-free or reference-based.
Reference-based [37] methods infer data membership of a given input text by comparing the target model’s outputs with the output of reference models, which were usually trained on the same data distribution. An example of such a method is LiRA [26] which estimates the likelihood ratio of a target example’s loss under models trained with and without the example, using Gaussian distributions to simplify the computation.
In contrast, reference-free methods aim to infer membership by applying different calculation methods on the probabilities of the token predictions. Among these methods is the LOSS Attack [31], which uses the loss values of a model; in the context of language models, this corresponds to the perplexity of the text. Perplexity is a measurement of how well a probability model predicts a sample, and it is calculated as the exponentiation of the negative average log-likelihood per token:
$\text{Perplexity}(P)=\exp\left(-\frac{1}{N}\sum_{i=1}^{N}\log P(t_{i}\mid t_{1},\ldots,t_{i-1})\right)$
where $N$ is the number of tokens, and $P(t_{i}\mid t_{1},\ldots,t_{i-1})$ is the conditional probability of the $i$-th token given its preceding tokens. The attack relies on the assumption that a lower perplexity indicates that the text is more familiar to the model, suggesting it may be part of the training set. The Zlib attack, as presented in the paper of Carlini et al.  [32], infers membership based on calculating the ratio between the log-likelihood of the text’s and its Zlib compression length. Newer attacks, such as the Neighbor attack[34], use a different language model than the target model in order to change selected words from a given text to create new sentences, noted as neighbors’ sentences, and then calculate the ratio between the original text’s perplexity and the perplexity of its neighbors’ sentences. Although the Neighbor attack showed some success, its computational cost is very high compared to other known methods. Attacks that outperform the Neighbor attack, with low computational requirements, are the Min-K[38], and Min-K%++ [39] methods which focus on the least confident predictions from the model’s output. Min-Kcalculates the average of the lowest kprobabilities from the model’s output, and Min-K%++ extends this by normalizing token log probabilities using the mean and variance. Lastly, two recently published attacks are RECALL [40], which measures the relative change in log-likelihood when conditioning the target text on non-member prefixes, and DC-PDD [41], which calibrates token probabilities using divergence from a reference corpus, effectively mitigating the impact of high-frequency tokens. Although each of the above attacks has been shown to be successful on some datasets and models, their performance was inconsistent across the examined datasets and models in each study. However, recent research performed by Maini et al. [24] showed that aggregating the results of multiple MIAs improves the accuracy of dataset membership inference. Although this is promising, based on their findings, we conclude that the success of this aggregated approach and its use in real-world scenarios relies on improved MIAs, and specifically, attacks capable of obtaining AUC results greater than 50%.
Common Models for evaluating MIAs on LLMs include open-source models with known pretraining data. Among these models is LLaMa 1 [7], created by Meta, which used the Pile dataset training data according to Meta’s publications. Recent research has also focused on the open-source series of models known as Pythia [42]. This suite of eight LLMs, ranging in size from 70M to 12B parameters, was trained on data including the Pile dataset, with all models processing the public data in the same order during their training phase.
3 Method
We introduce Tag&Tab, a novel method for identifying data used to pre-train LLMs, which is designed to be both resource- and time-efficient. Tag&Tab leverages common NLP techniques to tag keywords in the pretraining data and predict their occurrence using the target LLM.
Our method is based on the ability to strategically select words from the input that should be challenging for the LLM to predict. Successful prediction may imply that the model previously learned the input’s content during the pretraining process. In Tag&Tab, words are selected according to their entropy values. A high entropy value indicates that a word is less common in the input text compared to other words. Knowing that LLMs tend to memorize certain parts of the pretraining data that are rare or unique, the high-entropy words are more likely to be memorized [29], particularly in the context of their preceding words. Therefore, the LLM’s predictions for these words, given the preceding context, are likely to yield higher probabilities if the model has been trained on them, compared to other high-entropy words appearing in a different context that the model was not exposed to during the training.
Tag&Tab selects $K$ words with the highest entropy value in a sentence, which are referred to as Keywords. By selecting a small number of informative keywords, we aim to capture the semantic importance of the pretraining content. This approach hinges on a hypothesis that only $K$ selected keywords are needed to accurately predict the membership of the entire input text in the pretraining data, minimizing noise from other words in the input text [38].
Tag&Tab operates under black-box constraints regarding the target model, meaning that we can produce output from a given input text and observe the token probability of the model’s predictions, but we do not have access to the model’s weights [43, 25, 34]. Additionally, we assume that prior knowledge about part of the pretraining data or its distribution is known [7, 42, 44].
The Tag&Tab method consists of four stages, which are illustrated in Figure 1:
  1. 1.
    Preprocessing - First, the word entropy map is constructed using the Python package wordfreq [45], which provides frequency estimates for words in a specified language. The entropy for each word is calculated using the formula:
    $E(w_{i})=p(w_{i})\cdot\log_{2}p(w_{i})$
    In the preprocessing stage, the text is also split into individual sentences, using segmentation tools (e.g., the NLTK package [46]). To avoid selecting less informative keywords due to insufficient sentence length, sentences with fewer than a specified number of words are filtered out.
  2. 2.
    Tagging - From each sentence $S$ in the text file $T$, $K$ keywords are selected, targeting named entities and words with high entropy values. This selection is based on the word entropy map for high-entropy words, and named entities are identified during this process. Tools like spaCy [47] can be used to identify named entities.
  3. 3.
    Tabbing - This stage mimics the auto-completion feature found in interfaces like a command line, where it predicts and fills in the rest of the command based on the context of the preceding input. Using the target model $M$, we infer the log-likelihood of the entire text, and then the method calculates only the previously identified keywords. For each sentence $S\in T$ consisting of $n$ words $w_{1},w_{2},\ldots,w_{n}$, where each word $w_{i}$ is decomposed into tokens, denoted as $w_{i}=t_{i_{1}},t_{i_{2}},\ldots,t_{i_{m}}$,token $t_{i_{j}}$, given its preceding tokens, is calculated as $\log p_{M}(t_{i_{j}}|t_{i_{1}},\ldots,t_{i_{j-1}})$. We define the log-likelihood of a word $w_{i}$ using the log-likelihood of its first token $t_{i_{1}}$ given its preceding tokens, expressed as $\log p_{M}(t_{i_{1}}|t_{1_{1}},t_{1_{2}},\ldots,t_{i-1_{j}})$. As a result, we obtain:
    $\log p_{M}(w_{i}\mid\cdot)=\log p_{M}(t_{i_{1}}\mid\cdot)$
    The method selects the $K$ keywords from $S$ and computes the average log-likelihood of the keywords:
[…]
  4. 4.
    Inferring - In this stage, the method calculates the average of the keywords’ probability of all the sentences in the text $T$ and compares this value to a predetermined threshold $\gamma$, to determine the text’s membership.
4 Evaluation
In this section, we provide a detailed evaluation to showcase the effectiveness of Tag&Tab.
4.1 Model Comparison
To evaluate Tag&Tab in comparison to other reference-free baseline detection methods, we examined various open-source LLMs, including LLaMA 1 (7B, 13B, 30B) [7], Pythia (160M, 1.4B, 2.8B, 6.9B, 12B) [42]. Additionally, we used GPT-3.5 Turbo [^1] https://platform.openai.com/docs/models/gpt-3-5-turbo (trained on data up to September 2021222https://learn.microsoft.com/en-us/azure/ai-services/openai/concepts/models), given the partial knowledge that it was trained on known books, as discussed in previous studies [38, 44]. Models like LLaMA 1 and Pythia are well-suited for MIA evaluation with their transparency regarding pretraining datasets. This contrasts with newer models such as LLaMA 2 and 3, which lack such transparency.
4.2 Dataset Comparison
The experiments were conducted on the BookMIA333www.huggingface.co/datasets/swj0419/BookMIA [38], The Pile444www.huggingface.co/datasets/monology/pile-uncopyrighted/viewer/default/validation [12], and MIMIR555www.huggingface.co/datasets/iamgroot42/mimir [27] datasets, meeting the requirement that MIA evaluation datasets should be as comprehensive and diverse as possible [27], covering a wide range of text types while maintaining consistency in the distribution of text between the training and test sets.
The Pile is a collection of diverse textual sources designed to train and evaluate LLMs using open-source data, from which we utilized 10,000 samples each for training and testing from each domain. BookMIA focuses on evaluating MIAs with books known to have been memorized by GPT-3.5 Turbo, alongside newly published books from 2023, using 5,000 samples each for the member and non-member sets. It is important to note that 34 of the 50 ’member’ books in BookMIA overlap with the Gutenberg dataset, which is also known to have been part of the training corpora for models such as Pythia and LLaMA 1. MIMIR is a dataset built from The Pile, designed to evaluate memorization in LLMs. It contains data known to have been used in training across all Pythia model sizes, offering a unified benchmark for assessing membership inference. In our evaluation, we utilized around 2,000 samples per domain, combining approximately 1,000 ’member’ and 1,000 ’non-member’ samples.
It is important to note that we opted not to evaluate our method on the WikiMIA dataset666www.huggingface.co/datasets/swj0419/WikiMIA [38], as recent publications (e.g., [24]) questioned the reliability of the data, due to temporal shifts in writing styles and an insufficient number of samples.
4.3 Evaluation Approach
To assess our method’s performance, we followed a systematic process that begins with the input dataset. Each text file in the dataset is processed, with every sentence truncated to a maximum of 2,048 tokens to ensure a consistent input size. Sentences containing fewer than seven words are filtered out. We identify and save the top-K keywords for every sentence based on their entropy values. We documented the outcomes of selecting 1 to 10 keywords per sentence.
Once the top-k keywords are selected, the entire text is fed into the target model, which outputs the probability distribution for each token. Then we average the log-likelihoods of all tokens in the keywords, conditioned on their preceding tokens, assessing the model’s familiarity with the entire keyword.
Following the evaluation process suggested by Carlini et al.[26], we set a threshold value to determine attack performance, focusing on TPR at low FPR, denoted as T@F.
In addition, we report the area under the ROC curve (AUC score), which enhances understanding of the detection. The AUC score quantifies the overall performance of a classification method by considering TPRs and FPRs at all classification thresholds. Since AUC offers a comprehensive, threshold-independent evaluation metric, we do not need to determine a specific threshold $\gamma$ for our method.
4.4 Comparison with Baseline Methods
To benchmark Tag&Tab’s performance, we compared it to the performance of SOTA reference-free methods for detecting pretraining data. The baseline methods included LOSS Attack, Zlib Attack, Neighbor Attack, Min-KProb, Min-K%++ Prob, Max-K%Prob, RECALL, DC-PDD. Comprehensive descriptions of these attacks can be found in the Related Work section.
5 Results
In this section, we present two distinct case studies that utilize Tag&Tab, each exploring a different aspect of pretraining data detection in LLMs. Each case study is evaluated using the following metrics: AUC and TPR at a low FPR of 5% (T@F=5%).
Method LLaMa-7b LLaMa-13b LLaMa-30b Pythia-6.9b Pythia-12b GPT-3.5 Average
AUC T@F5 AUC T@F5 AUC T@F5 AUC T@F5 AUC T@F5 AUC T@F5 AUC T@F5
Neighbor 64.7 26.8 71.2 38.3 89.6 72.8 64.9 25.8 70.9 36.1 96.3 87.5 76.3 47.9
Loss 58.9 24.6 69.6 43.1 88.7 74.3 62.4 23.7 68.5 32.2 96.6 89.7 74.1 47.9
Zlib 52.5 21.9 66.8 41.6 88.7 74.1 54.6 18.7 61.1 25.4 95.6 88.1 69.9 44.9
Min-20.0% Prob 61.1 24.3 70.3 42.4 86.7 70.0 65.2 24.6 70.0 34.0 95.1 86.2 74.7 46.9
MinK++-20.0% Prob 59.7 23.0 67.7 37.6 78.0 60.2 59.1 19.6 56.2 20.1 95.1 86.2 69.3 41.1
Max-20.0% Prob 51.4 15.3 65.6 33.9 87.4 69.3 50.6 12.7 58.8 20.2 96.0 90.6 68.3 40.3
ReCaLL 57.8 22.3 69.8 41.9 84.1 64.0 66.2 29.1 71.7 36.8 73.7 50.3 70.5 41.1
DC-PDD 60.7 26.9 71.0 46.9 88.3 76.7 67.9 33.5 73.6 44.2 95.2 88.7 76.1 52.8
Ours (Tag&Tab K=4) 69.1 28.2 77.8 47.5 91.4 76.4 71.7 29.6 75.4 35.9 97.0 89.6 80.4 51.2
Ours (Tag&Tab K=10) 67.0 26.1 76.6 46.2 91.4 76.5 71.9 30.2 75.6 35.6 95.9 87.3 79.7 50.3
Table 1: Detection of data from the BookMIA dataset used in pretraining seven models using Tag&Tab and six baseline MIAs, evaluated in terms of the AUC and T@F=5%. The last two rows compare the Tag&Tab method when selecting four and ten keywords. The best results are highlighted in bold, and the second-best results are underscored.
5.1 Case Study 1: Detecting Specific Pretraining Data in LLMs
This case study focuses on the precision of detecting specific pretraining data in LLMs. To do so, we designed a well-engineered attack aimed at inferring whether copyrighted data was part of the pretraining of the examined LLMs. In contrast to Case Study 2, the data used for training the target model and the data identified as non-members here do not originate from the same database. Using the BookMIA dataset, we simulate the possibility of having knowledge of part of a model’s pretraining data, which can be used to infer specific text files suspected of being included in the pretraining data of the target model. For validation, non-members were collected from books that were released after the publication of the target model, ensuring that these texts could not be part of its pretraining data.
To determine the optimal number of keywords to select, we evaluate the results by selecting between 1 and 10 keywords from each sentence The results shown in Figure 2 demonstrate that the optimal number of keywords required to ensure effective detection depends on the model architecture. As observed, for different sizes of the LLaMa 1 models, the optimal number of keywords ranged from 2 to 3, while for the Pythia models and GPT-3.5 turbo, the optimal number of tagged keywords was 7. On average the best results across all models were obtained when $K=4$. resulting in an average AUC score of 80.4%.
Table 1 summarizes the results of Tag&Tab and reference-free baseline attacks.
Tag&Tab outperforms all other attacks in terms of the AUC score, achieving an average improvement in the AUC compared to the other attacks, ranging from 4.1%-12.1% for Tag&Tab K=4. The performance of Tag&Tab K=4 surpasses that of other attacks in terms of the T@F=5% on LLaMa-7b and LLaMa-13b models. However, for most other models, the DC-DPP attack achieves better results, with Tag&Tab consistently ranking second in performance. As the size of the tested LLM increases, the AUC scores of the MIAs also increase due to the model’s memorization capacity [29]. This can be seen in the results obtained for Tag&Tab which obtained very high AUC scores: (1) 91.4% on LLaMa-30b compared to 69.1% on LLaMa-7b, (2) 75.4% on Pythia-12b compared to 71.7% Pythia-6.9b, and (3) 97% on GPT-3.5 Turbo, which is the largest model tested.
Finally, we also investigated the impact of our method’s word selection stage, tagging, in which the highest $K$ entropy words are selected, compared to a random selection of tokens from the sentence. The results indicate that selecting the highest $K$ entropy words significantly improves performance across all models, achieving an average AUC score of 80.4%, compared to an average AUC score of 64.4% with a random selection of $K$ tokens.
Figure 2: AUC scores as a function of the number of tagged keywords for the examined models on the BookMIA dataset. Yellow dots indicate optimal performance: 2-3 keywords for LLaMa 1, 7 for Pythia and GPT-3.5 turbo, and 4 on average (AVG).
Table 2: Comparison of AUC results for our method, Tag&Tab, against baseline methods on the MIMIR and The PILE benchmarks. The upper table presents the best results from Tag&Tab and baseline methods on four datasets from the MIMIR benchmark, while the lower table shows the best results on four datasets from The PILE benchmark. The best results for each dataset and model size are highlighted in bold, and the second-best AUC is underscored.
Method DM Mathematics Github Pile CC C4
160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B
Loss 0.85 0.76 0.84 0.68 0.86 0.80 0.85 0.86 0.88 0.88 0.53 0.54 0.54 0.55 0.55 0.50 0.51 0.51 0.51 0.51
Zlib 0.68 0.59 0.66 0.55 0.69 0.84 0.88 0.89 0.90 0.90 0.51 0.53 0.53 0.54 0.54 0.51 0.51 0.51 0.51 0.51
Min-20% Prob 0.61 0.53 0.70 0.50 0.82 0.80 0.85 0.86 0.88 0.88 0.52 0.53 0.54 0.55 0.55 0.51 0.51 0.51 0.51 0.50
Max-20% Prob 0.63 0.67 0.61 0.58 0.51 0.78 0.85 0.85 0.87 0.86 0.52 0.53 0.53 0.53 0.54 0.51 0.50 0.50 0.50 0.50
Min-K++-20% Prob 0.81 0.79 0.66 0.81 0.73 0.57 0.57 0.61 0.63 0.66 0.51 0.50 0.52 0.53 0.53 0.52 0.51 0.51 0.50 0.50
RECALL 0.80 0.73 0.78 0.64 0.86 0.79 0.76 0.74 0.71 0.72 0.53 0.54 0.54 0.55 0.55 0.51 0.51 0.51 0.51 0.51
DC-PDD 0.90 0.86 0.86 0.85 0.86 0.87 0.91 0.92 0.93 0.93 0.54 0.55 0.56 0.57 0.57 0.51 0.51 0.51 0.51 0.51
Ours (Tag&Tab K=4) 0.96 0.96 0.96 0.95 0.95 0.78 0.82 0.83 0.84 0.85 0.54 0.56 0.56 0.57 0.57 0.53 0.52 0.52 0.52 0.51
Ours (Tag&Tab K=10) 0.92 0.92 0.93 0.92 0.95 0.79 0.83 0.84 0.85 0.86 0.55 0.56 0.56 0.57 0.56 0.53 0.52 0.52 0.52 0.51
Method Ubuntu IRC Gutenberg EuroParl Average
160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B
Loss 0.63 0.59 0.60 0.58 0.58 0.53 0.53 0.53 0.53 0.53 0.52 0.52 0.50 0.52 0.51 0.67 0.67 0.69 0.66 0.70
Zlib 0.52 0.52 0.53 0.54 0.54 0.53 0.60 0.53 0.53 0.53 0.51 0.51 0.50 0.51 0.51 0.63 0.63 0.65 0.62 0.66
Min-20% Prob 0.58 0.57 0.52 0.51 0.52 0.53 0.53 0.53 0.53 0.60 0.53 0.54 0.52 0.50 0.51 0.61 0.61 0.65 0.61 0.69
Max-20% Prob 0.69 0.69 0.71 0.68 0.67 0.67 0.73 0.60 0.67 0.67 0.53 0.54 0.55 0.53 0.55 0.61 0.64 0.62 0.62 0.60
Min-K++-20% Prob 0.52 0.51 0.52 0.54 0.61 0.67 0.60 0.60 0.60 0.60 0.54 0.53 0.51 0.51 0.51 0.60 0.59 0.57 0.62 0.61
RECALL 0.72 0.64 0.69 0.64 0.60 0.53 0.80 0.67 0.73 0.80 0.51 0.51 0.51 0.55 0.57 0.67 0.64 0.65 0.62 0.68
DC-PDD 0.58 0.53 0.53 0.53 0.53 0.53 0.60 0.60 0.53 0.53 0.51 0.52 0.50 0.51 0.54 0.70 0.70 0.72 0.71 0.70
Ours (Tag&Tab K=4) 0.64 0.65 0.64 0.66 0.64 0.67 0.67 0.67 0.67 0.67 0.55 0.54 0.55 0.54 0.56 0.70 0.72 0.73 0.72 0.73
Ours (Tag&Tab K=10) 0.61 0.63 0.62 0.61 0.62 0.60 0.67 0.67 0.67 0.67 0.56 0.54 0.55 0.54 0.55 0.70 0.71 0.73 0.72 0.72
5.2 Case Study 2: Detecting Various Types of Pretraining Data in LLMs
This case study examines the Tag&Tab effectiveness and robustness in detecting various types of pretraining data in LLMs. We evaluate our method on various sizes of the Pythia model and compare its effectiveness against baseline attacks on 7 different types of textual data in the Pile dataset. Table 2 summarizes the results obtained when targeting five Pythia model sizes ranging from 160M to 12B parameters, tested with two configurations of tagged keywords: 4 and 10.
Tag&Tab ($K=4$) consistently outperforms all baseline methods on average across the evaluated models, establishing itself as the most effective approach overall. Tag&Tab ($K=10$) ranks as the second-best method, demonstrating strong performance but falling short of the results achieved with $K=4$. Notably, Tag&Tab ($K=4$) achieves either the best or second-best results in the majority of tested types of textual daat. Even when it does not lead, its performance remains competitive, providing a robust alternative to the leading method.
We can also see that while Tag&Tab ($K=10$) selects more keywords, thus increasing the number of probabilities considered for text membership inference, the results are consistently lower compared to Tag&Tab when $K=4$. This supports the hypothesis that selecting a smaller number of keywords allows the method to extract noise-free information from the model.
Despite outperforming baseline membership inference attacks (MIAs), the AUC achieved by Tag&Tab can still be relatively low in certain cases, hovering around 55%. However, new research performed by Maini et al. [24] shows that aggregating multiple MIAs improves dataset membership inference accuracy, motivating the need for better attacks that achieve AUC over 50% for improved aggregated attack performance.
6 Conclusion
We present Tag&Tab, a novel black-box method for detecting pretraining data in LLMs. The method’s focus on the semantic and contextual relevance of words in the text enhances its detection capabilities. Tag&Tab was shown to outperform SOTA attacks and demonstrates consistently high performance across diverse types of textual data. In our comprehensive evaluation on 8 different types of textual data from three datasets (the Pile, MIMIR, and BookMIA datasets) and six different LLMs (LLaMa 1-7b, 13b, 30b, Pythia-160m, 1.4b, 2.4b, 6.9b, 12b, GPT-3.5 Turbo) of varying sizes and architectures. Our study confirms that the selection of high-entropy keywords improves membership inference attack results, further validating our approach. Future work could explore extending the Tag&Tab method by incorporating dimensions such as the context, and placement of keywords within the document. Additionally, developing new MIAs that leverage advanced NLP techniques to focus on the significance of words within the text could further enhance the ability to detect pertaining data in LLMs.